# Reversibility in the Seifert-fibered spaces

### v2

@@ Abstract.
An element $a$ in a group $\Gamma$ is called reversible if there exists $g\in\Gamma$ such that $gag^{-1}=a^{-1}$. The reversible elements are also known as ‘real elements’ or ‘reciprocal elements’ in literature. In this paper, we classify the reversible elements in Fuchsian groups, and use this classification to find all reversible elements in a Seifert-fibered group. In the last section we apply this classification to the braid groups, particularly to the braid group on $3$ strands.
@@ 1. Introduction
An element $a$ in a group $\Gamma$ is said to be ‘reversible’ if it is conjugate to its own inverse, and the conjugating elements are called ‘reverser’ or ‘reversing symmetry’. If the conjugating element is also an involution, then it is called ‘strongly reversible’ element. It has also been called by other terms like ‘reciprocals’ or ‘real elements’ in other literature, e.g., [BR], [La] etc. In this paper our discussion is limited to some properties of reversible elements, and a discussion of reversible elements in some discrete geometric groups. Reversibility has its own geometric motivations and has been discussed widely in literature. For instance, in a Fuchsian group, the collection of conjugacy classes of reversible elements have a 1-1 correspondence to the conjugacy classes of infinite dihedral groups contained in that Fuchsian group. But in general, supposing $G$ is a 3-manifold group (the fundamental group of some 3-manifold), if $G$ is torsion free then a reversible element $g$ in $G$ is represented as a loop which is homotopic to its own inverse (i.e., it will form an embedded Klein bottle with the reversing symmetry). So, it can be considered that finding a reversible element with a reverser in a 3-manifold is equivalent to finding an embedded Klein bottle upto homotopy inside the manifold. In this paper we are specially interested in Seifert fibered manifolds since the fundamental group of a Seifert fibered manifold is a universal extension of Fuchsian groups or non-orientable 2-orbifold groups.
@@
A Seifert fibered manifold is a compact $3$-manifold which can be seen as a $S^{1}$ bundle over a $2$-dimensional orbifold as the base space, such that every fiber has a solid fibered torus neighbourhood. These spaces were first studied by Herbert Seifert in his doctoral thesis [ST], and they form an important example of $3$-manifolds. In fact, all the the compact orientable $3$-manifolds in $6$ of the $8$ possible $3$-manifold geometric structures can be seen as Seifert fibered spaces. If we consider the space of fibers of $M$ with a quotient topology, we get a resulting surface called the orbit surface of $M$. We denote the orbit surface by $G$. The orbit surface can also be seen as a quotient of the base space of $M$. There is a map $\phi:\pi_{1}(G)\to\{-1,1\}$ called the classifying homomorphism. A loop $\alpha$ on $G$ has image $\phi(\alpha)=1$ if the fiber above $\alpha$ is mapped to itself in an orientation preserving way as we follow the loop $\alpha$, and has image $\phi(\alpha)=-1$ otherwise. The classifying homomorphism plays a crucial role in determining the fiber type of a Seifert fibered space. Following Seifert’s work and Brin’s note in [Br], the closed Seifert fibered spaces can be completely classified based on certain symbols. If $M$ is a closed and oriented Seifert fibered space, then $M$ is uniquely determined by a symbol
$(O,o,g|b,\mu_{1},\beta_{1},\dots,\mu_{n},\beta_{n})$
or
$(O,n,k|b,\mu_{1},\beta_{1},\dots,\mu_{n},\beta_{n}).$
Here $O$ refers to the orientability of $M$, $o$ and $n$ refer to the orientability or non-orientability of the base space respectively, $g$ is the genus of the orientable base space, and $k$ is the number of crosscaps of the non-orientable base space. $b$ is an integer depending on the fibration of the space, and the $\mu_{i}$ and $\beta_{i}$ depend on the fiber type of the $i^{th}$ essential fiber. If $M$ is non-orientable, then there are $4$ possible symbols associated:
${(N,o,g|b,\mu_{1},\beta_{1},\dots,\mu_{n},\beta_{n})}$
if the base space is orientable, and otherwise, one of the follwing $3$ types based on the classifying homomorphism of the base space:
$(N,n,\textrm{I},k|b,\mu_{1},\beta_{1},\dots,\mu_{n},\beta_{n});$
$(O,n,\textrm{II},k|b,\mu_{1},\beta_{1},\dots,\mu_{n},\beta_{n});$
$(O,n,\textrm{III},k|b,\mu_{1},\beta_{1},\dots,\mu_{n},\beta_{n}).$
@@
Given any such Seifert fibered space, this paper is concerned with checking the reversibility of the elements of $\pi_{1}(M)$, the fundamental group of $M$. The fundamental group of $M$ depends on the orientability of the orbit surface $G$. If $G$ is orientable, it can be shown that the fundamental group of $M$ is of the form
(1.1) $\pi_{1}(M)=$ $\langle a_{i},b_{i},c_{i},d_{i},h|a_{i}ha_{i}^{-1}=h^{\phi(a_{i})},$
(1.2) $b_{i}hb_{i}^{-1}=h^{\phi(b_{i})},c_{i}hc_{i}^{-1}=h,d_{i}hd_{i}^{-1}=h^{\phi(d_{i})},$
(1.3) $c_{i}^{\mu_{i}}=h^{\beta}_{i},\Pi[a_{i},b_{i}]\Pi c_{i}\Pi d_{i}h^{b}=1\rangle.$
@@
(1.4) $\pi_{1}(M)=$ $\langle x_{i},c_{i},d_{i},h|x_{i}hx_{i}^{-1}=h^{\phi(x_{i})},c_{i}hc_{i}^{-1}=h,d_{i}hd_{i}^{-1}=h^{\phi(d_{i})},$
(1.5) $c_{i}^{\mu_{i}}=h^{\beta_{i}},\Pi x_{i}^{2}\Pi c_{i}\Pi d_{i}h^{b}=1\rangle.$
@@
$\pi_{1}(M)/\langle h\rangle=$ $\langle a_{i},b_{i},c_{i},d_{i}|c_{i}^{\mu_{i}}=1,\Pi[a_{i},b_{i}]\Pi c_{i}\Pi d_{i}=1\rangle.$
@@
$\pi_{1}(M)/\langle h\rangle=$ $\langle x_{i},c_{i},d_{i}|c_{i}^{\mu_{i}}=1,\Pi x_{i}^{2}\Pi c_{i}\Pi d_{i}=1\rangle.$
@@
If $M$ has an orientable orbit surface and $\phi$ is nontrivial, the reversible elements of $\pi_{1}(M)$ are the elements conjugate to $h$ and $c_{i}^{\mu_{i}}$ for all $i$ and the conjugates of the elements of either of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$ with $\phi(k)=-1$ or of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{-\mu_{j}}{2}}k^{-1}$ with $\phi(k)=1$, where $\mu_{i}$ and $\mu_{j}$ are even, and $k$ is any element in $\pi_{1}(M)$, $\beta_{i}=\beta_{j}$. $\pi_{1}(M)$ has reversible elements as the elements conjugate to $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{-\mu_{j}}{2}}k^{-1}$ if $\phi$ is trivial.
We also similarly classify the reversible elements in case of Seifert fibered spaces with non-orientable orbit surfaces. In Section 2, we discuss about the reversible elements of fundamental groups of two-dimensional hyperbolic orbifolds, and we classify the reversible elements of such groups. This classification is then applied to the fundamental groups of any Seifert-fibered space in Section 3. Section 4 then discusses an application of the results of Section 3 on the fundamental group of $B_{3}$, the braid group on $3$ strands. The possibility of using the same methods for any $B_{n}/\langle h\rangle$ is also discussed there. After completion of the first draft, the authors were informed by Prof. Masakazu Teragaito regarding the article [HMT] written by Himeno, Motegi, and Teragaito, where they classify 3-manifolds whose fundamental groups admit generalised torsion elements of order two, and further classify such elements. Their proof is topological and completely different from the proof in this paper. The two articles were written completely independently and at around the same time. The proof in [HMT] crucially uses important results by Jaco-Shalen and Hass, while the techniques in this article are simpler by comparison.
@@ 2.1. Reversible classes of Fuchsian group
In [DG2, Remark 1], we have seen the classifications of the $z$-classes for hyperbolic elements. In the same way, we can see the classifications of reciprocal elements for any finite type Fuchsian groups. We have discussed the reciprocal elements of Hecke groups in [DG1], which is a particular class of Fuchsian groups. From now on, the hyperbolic axis for an hyperbolic element $h$ is written as $I_{h}$.
@@
Proof. $\Gamma$ is in $\rm PSL(2,\mathbb{R})$, and so we will conjugate $\Gamma$ by a Möbius map $f$ such that in the new group after conjugation, $h$ will be a hyperbolic element in the Poincaré disk where the hyperbolic axis of $h$ passes through $(0,0)$. Now, the reciprocator for $h$ is $g$, which is an involution having fixed point $z$. So, let us suppose $z$ does not lie on $I_{h}$ and $I_{h}$ is a straight line passing through the origin. Then, $z$ lies one side of the axis. In our earlier work [DG1, Lemma 2.2] we have shown that in the Fuchsian group, the reciprocators of a reciprocal element fix the hyperbolic axis of that particular hyperbolic element and interchange the fixed points of hyperbolic axis. Now, we apply the map $g$ to the disk. This map is a $\pi$-rotation along $I_{h}$ of the disk, and therefore the point $z$ must change its position. This is contradiction. So, the fixed point of $g$ must lie on $I_{h}$. ∎
@@
Proof. We first show that if $g$ is reciprocal, then $I_{g}$ passes through at least one even order vertex in the fundamental domain. First, the fixed points of the reciprocator of $g$ will always be a vertex of the domain. From [DG2, Theorem 3], we obtain that $I_{g}$ passing through a fixed domain represents a $z$-class of that Fuchsian group. And if it is reciprocal then there is a fixed point of one of the reciprocator of $g$ that lies on $I_{g}$ by 2.2, and it is an even ordered-vertex of the fundamental domain (from the definition of fundamental domain). This proves one side of the statement. Conversely, if the axis of the primitive hyperbolic element $g$ passes through an even-order vertex $z$ of the tiling which is done by the fundamental domain, then let $k$ be the elliptic finite order element in the Fuchsian group such that, $k(z)=z$ and $k^{2m}=Id$. So, $k^{m}$ is an involution and it is a is $\pi$-rotation. Then, it will interchange the fixed points of $g$, and so $k^{m}gk^{m}=g^{-1}$ since $g$ is primitive. This shows that any non-trivial power of $g$ is reciprocal. This proves the lemma. ∎
@@ 2.3. Reversible classes of non-orientable surface fundamental groups
If $k>2$, there is no finite order element in the fundamental group. Then our claim is that there are no more reversible elements in $S$. If on the contrary there exists any loop $l$ that homotopes to its own inverse, then that element along with the reverser will form a Klein bottle which is a closed surface (if the reverser is infinite order). That implies, there is an embedding of the Klein bottle in $S$. Since both $S$ and the Klein bottle have same dimension and are both complete surfaces, therefore by the Hopf–Rinow theorem, they must be homeomorphic. This contradicts our hypothesis and hence proves the lemma.
@@
Infinite order reversible elements in Γsuperscriptnormal-$\Gamma^{\prime}$ are strongly reversible elements.
The proof is similar to the proof of [DG1, Lemma 3.4 ].
Proof. The group $\Gamma^{\prime}$ is a discrete subgroup of $\rm PGL(2,\mathbb{R})$. Let $r$ in ${\bf H}^{2}/\Gamma^{\prime}$ be a reversible element in $\Gamma^{\prime}$ Then there is an element $b$ in $\Gamma^{\prime}$ such that $b^{-1}rb=r^{-1}$. At first we consider that $r$ is orientation preserving isometry in ${\bf H}^{2}$ (which implies that will not be a parabolic element). Then, $b$ either fixes the fixed points of $r$ or reflects the fixed points to each other. If $b$ is fixes the fixed points of $r$ then it will either commute with $r$ or reverse $r$. In the case where it commutes with $r$ it leads us to a contradiction that $b$ is not a reverser, and in the case where it is reversing, $b$ will be an involution. If $b$ interchanges the fixed points, then $b$ must be an involution. In $r$ is a orientation reversing element, then $r^{2}$ will be orientation preserving element with the same reverser $b$. From the previous case, $b^{2}=1$. Hence, in any case, $r$ will be strongly reversible. ∎
@@
The infinite ordered reversible elements in Γsuperscriptnormal-$\Gamma^{\prime}$ are conjugate to $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$ where $\mu_{i}$ and $\mu_{j}$ are even and $k\in\Gamma^{\prime}$.
Proof. Since there is no finite order element in $\Gamma^{\prime}$ other than the $c_{i}$, this theorem follows from the previous lemma. ∎

### v3

# Reversible and other generalised torsion elements in Seifert-fibered spaces

@@ Abstract.
An element $a$ in a group $\Gamma$ is called reversible if there exists $g\in\Gamma$ such that $gag^{-1}=a^{-1}$. The reversible elements are also known as ‘real elements’ or ‘reciprocal elements’ in literature. In this paper, we classify the reversible elements in Fuchsian groups, and use this classification to find all reversible elements in a Seifert-fibered group. We then apply the classification to the braid groups, particularly to the braid group on $3$ strands. We further study generalised 3-torsion elements in PSL(2,Z), and use this to analyse the existence of generalised 3-torsion elements in Seifert-fibered spaces in general, and braid groups on 3 strands in particular.
@@ 1. Introduction
A non-trivial element $a$ in a group $\Gamma$ is said to be ‘reversible’ if it is conjugate to its own inverse, and the conjugating elements are called ‘reverser’ or ‘reversing symmetry’. If the conjugating element is also an involution, then it is called ‘strongly reversible’ element. It has also been called by other terms like ‘reciprocals’ or ‘real elements’ in other works, e.g., [BR], [La] etc. In this paper our scope is limited to some properties of reversible elements, and the existence of reversible elements in some discrete groups. Reversibility has its own geometric motivations and has been studied widely in literature. For instance, in a Fuchsian group, the collection of conjugacy classes of reversible elements have a 1-1 correspondence to the conjugacy classes of infinite dihedral groups contained in that Fuchsian group. But in general, supposing $G$ is a 3-manifold group (the fundamental group of some 3-manifold), if $G$ is torsion free then a reversible element $g$ in $G$ is represented as a loop which is homotopic to its own inverse (i.e., it will form an embedded Klein bottle with the reversing symmetry). So, finding a reversible element with a reverser in a 3-manifold is equivalent to finding an embedded Klein bottle up to homotopy inside the manifold. In this paper we are specially interested in Seifert fibered manifolds since the fundamental group of a Seifert fibered manifold is a universal extension of Fuchsian groups or non-orientable 2-orbifold groups.
Furthermore, an element $g$ in a group $G$ is said to be a generalised $n$-torsion element if it satisfies the condition
$gh_{1}gh_{1}^{-1}h_{2}gh_{2}^{-1}\dots h_{n-1}gh^{-1}_{n-1}=Id.$
Generalised $n$-torsion elements are conjugacy invariant. Reversible elements are therefore also known as generalised $2$-torsion elements ( see [HMT]). We also study the existence of generalised $3$-torsion elements in $\rm PSL(2,\mathbb{Z})$ and use it to find such elements in Seifert-fibered spaces.
@@
A Seifert fibered manifold is a compact $3$-manifold which can be seen as a $S^{1}$ bundle over a $2$-dimensional orbifold as the base space, such that every fiber has a solid fibered torus neighbourhood. These spaces were first studied by Herbert Seifert in his doctoral thesis [ST], and they form an important example of $3$-manifolds. In fact, all the the compact orientable $3$-manifolds in $6$ of the $8$ possible $3$-manifold geometric structures can be seen as Seifert fibered spaces. Seifert fibered spaces have been widely studied, for instance in [JS], [LR], [KR]. If we consider the space of fibers of a Seifert fibered space $M$ with a quotient topology, we get a resulting surface called the orbit surface of $M$. We denote the orbit surface by $G$. The orbit surface can also be seen as a quotient of the base space of $M$. There is a map $\phi:\pi_{1}(G)\to\{-1,1\}$ called the classifying homomorphism. A loop $\alpha$ on $G$ has image $\phi(\alpha)=1$ if the fiber above $\alpha$ is mapped to itself in an orientation preserving way as we follow the loop $\alpha$, and has image $\phi(\alpha)=-1$ otherwise. The classifying homomorphism plays a crucial role in determining the fiber type of a Seifert fibered space.
@@
Given any such Seifert fibered space, this paper is concerned with checking the reversibility of the elements of $\pi_{1}(M)$, the fundamental group of $M$. The fundamental group of $M$ depends on the orientability of the orbit surface $G$. If $G$ is orientable, it can be shown following Seifert’s work and Brin’s note in [Br], that the fundamental group of $M$ is of the form
(1.1) $\pi_{1}(M)=$ $\langle a_{i},b_{i},c_{i},d_{i},h~{}|~{}a_{i}ha_{i}^{-1}=h^{\phi(a_{i})},$
$b_{i}hb_{i}^{-1}=h^{\phi(b_{i})},c_{i}hc_{i}^{-1}=h,d_{i}hd_{i}^{-1}=h^{\phi(d_{i})},$
$c_{i}^{\mu_{i}}=h^{\beta}_{i},\prod~{}[a_{i},b_{i}]~{}\prod c_{i}~{}\prod d_{i}h^{b}=1\rangle.$
@@
(1.2) $\pi_{1}(M)=\langle x_{i},c_{i},d_{i},h~{}|~{}x_{i}hx_{i}^{-1}=h^{\phi(x_{i})},$
$c_{i}hc_{i}^{-1}=h,d_{i}hd_{i}^{-1}=h^{\phi(d_{i})},c_{i}^{\mu_{i}}=h^{\beta_{i}},\prod x_{i}^{2}~{}\prod c_{i}\prod d_{i}h^{b}=1\rangle.$
@@
$\pi_{1}(M)/\langle h\rangle=$ $\langle a_{i},b_{i},c_{i},d_{i}|c_{i}^{\mu_{i}}=1,\prod[a_{i},b_{i}]~{}\prod c_{i}~{}\prod d_{i}=1\rangle.$
@@
$\pi_{1}(M)/\langle h\rangle=$ $\langle x_{i},c_{i},d_{i}~{}|~{}c_{i}^{\mu_{i}}=1,\prod~{}x_{i}^{2}\prod~{}c_{i}\prod~{}d_{i}=1\rangle.$
@@
If a Seifert fibered space $M$ has an orientable orbit surface. If $\phi$ is nontrivial, the reversible elements of $\pi_{1}(M)$ are the following:
  1. (1)
    the elements conjugate to any power of $h$.
  2. (2)
    the conjugates of the elements of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$ with $\phi(k)=-1$ where $\mu_{i}$ and $\mu_{j}$ are even and $k$ is any element in $\pi_{1}(M)$, and $\beta_{i}=\beta_{j}$.
  3. (3)
    the conjugates of the elements of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{-\mu_{j}}{2}}k^{-1}$ with $\phi(k)=1$, where $\mu_{i}$ and $\mu_{j}$ are even and $k$ is any element in $\pi_{1}(M)$, and $\beta_{i}=\beta_{j}$.
If $\phi$ is trivial, $\pi_{1}(M)$ has reversible elements as the elements conjugate to $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{-\mu_{j}}{2}}k^{-1}$.
We also similarly classify the reversible elements in case of Seifert fibered spaces with non-orientable orbit surfaces.
In Section 2, we study the fundamental groups of two-dimensional orbifolds, and classify the reversible elements of such groups. This classification is then applied to the fundamental groups of any Seifert-fibered space in Section 3. Section 4 then discusses an application of the results of Section 3 on the fundamental group of $B_{3}$, the braid group on $3$ strands. The possibility of using the same methods for any $B_{n}/\langle h\rangle$ is also discussed there. In Section 5, we study the existence of generalised $3$-torsion elements in Seifert-fibered spaces, and apply it in $\pi_{1}(B_{3})$ to find such elements. We also find conditions for the existence of generalised $n$-torsion elements.
After completion of the first draft, the authors were informed by Prof. Masakazu Teragaito regarding the article [HMT] written by Himeno, Motegi, and Teragaito, where they classify 3-manifolds whose fundamental groups admit generalised torsion elements of order two, and further classify such elements. Their proof is topological and completely different from the proof in the Section 3 of this paper. The two articles were written completely independently and at around the same time. The proof in [HMT] crucially uses important results by Jaco-Shalen and Hass, while the techniques in this article are simpler in comparison.
@@ 2.1. Reversible classes of Fuchsian group
We can see the classifications of reciprocal elements for any finite type Fuchsian groups. We have discussed the reciprocal elements of Hecke groups in [DG], which is a particular class of Fuchsian groups.
From now on, the hyperbolic axis for an hyperbolic element $h$ is written as $I_{h}$.
@@
Proof. $\Gamma$ is in $\rm PSL(2,\mathbb{R})$, and so we will conjugate $\Gamma$ by a Möbius map $f$ such that in the new group after conjugation, $h$ will be a hyperbolic element in the Poincaré disk where the hyperbolic axis of $h$ passes through $(0,0)$. Now, the reciprocator for $h$ is $g$, which is an involution having fixed point $z$. So, let us suppose $z$ does not lie on $I_{h}$ and $I_{h}$ is a straight line passing through the origin. Then, $z$ lies one side of the axis. In our earlier work [DG, Lemma 2.2] we have shown that in the Fuchsian group, the reciprocators of a reciprocal element fix the hyperbolic axis of that particular hyperbolic element and interchange the fixed points of hyperbolic axis. Now, we apply the map $g$ to the disk. This map is a $\pi$-rotation along $I_{h}$ of the disk, and therefore the point $z$ must change its position. This is contradiction. So, the fixed point of $g$ must lie on $I_{h}$. ∎
@@
Proof. We first show that if $g$ is reciprocal, then $I_{g}$ passes through at least one even order vertex in the fundamental domain. First, the fixed points of the reciprocator of $g$ will always be a vertex of the domain. We know that $I_{g}$ passing through a fixed domain represents a $z$-class of that Fuchsian group. And if it is reciprocal then there is a fixed point of one of the reciprocator of $g$ that lies on $I_{g}$ by 2.2, and it is an even ordered-vertex of the fundamental domain (from the definition of fundamental domain). This proves one side of the statement.
Conversely, if the axis of the primitive hyperbolic element $g$ passes through an even-order vertex $z$ of the tiling which is done by the fundamental domain, then let $k$ be the elliptic finite order element in the Fuchsian group such that, $k(z)=z$ and $k^{2m}=Id$. So, $k^{m}$ is an involution and it is a is $\pi$-rotation. Then, it will interchange the fixed points of $g$, and so $k^{m}gk^{m}=g^{-1}$ since $g$ is primitive. This shows that any non-trivial power of $g$ is reciprocal. This proves the lemma.
@@ 2.3. Reversible classes of non-orientable surface fundamental groups
If $k>2$, there is no finite order element in the fundamental group. Then our claim is that there are no more reversible elements in $S$. If on the contrary there exists any loop $l$ that homotopes to its own inverse, then $\pi_{1}(S)$ will no longer be a biorderable group. This will lead us to a contradiction. Hence, this proves the lemma.
@@
Infinite order reversible elements in $\Gamma^{\prime}$ are strongly reversible elements, and every reverser is an involution.
The proof is similar to the proof of [DG, Lemma 3.4 ].
Proof. The group $\Gamma^{\prime}$ is a discrete subgroup of $\rm PGL(2,\mathbb{R})$. Let $r$ in ${\bf H}^{2}/\Gamma^{\prime}$ be a reversible element in $\Gamma^{\prime}$ Then there is an element $b$ in $\Gamma^{\prime}$ such that $b^{-1}rb=r^{-1}$. At first we consider that $r$ is orientation preserving isometry in ${\bf H}^{2}$ (which implies that will not be a parabolic element). Then, $b$ either fixes the fixed points of $r$ or reflects the fixed points to each other. If $b$ fixes the set of fixed points of $r$ then it will either commute with $r$ or reverse $r$. In the case where it commutes with $r$ it leads us to a contradiction that $b$ is not a reverser, and in the case where it is reversing, $b$ will be an involution. If $b$ interchanges the fixed points, then $b$ must be an involution. So, every reverser is an involution in $\Gamma^{\prime}$
In $r$ is a orientation reversing element, then $r^{2}$ will be orientation preserving element with the same reverser $b$. From the previous case, $b^{2}=1$. Hence, in any case, $r$ will be strongly reversible. ∎
@@
The infinite order reversible elements in $\Gamma^{\prime}$ are conjugate to $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$ where $\mu_{i}$ and $\mu_{j}$ are even and $k\in\Gamma^{\prime}$.
Proof. Since there is no finite order element in $\Gamma^{\prime}$ other than the elements conjugate to $c_{i}$, this theorem follows from the previous lemma. ∎